Target Costing

Laser Impressions, Inc., manufactures color laser printers. Model J20 presently sells for $350 and has a total product cost of $280, as follows:

Direct materials $200
Direct labor 50
Factory overhead 30
Total $280

It is estimated that the competitive selling price for color laser printers of this type will drop to $330 next year. Laser Impressions has established a target cost to maintain its historical markup percentage on product cost. Engineers have provided the following cost-reduction ideas:

  1. Purchase a plastic printer cover with snap-on assembly, rather than with screws. This will reduce the amount of direct labor by 9 minutes per unit.
  2. Add an inspection step that will add six minutes per unit of direct labor but reduce the materials cost by $7 per unit.
  3. Decrease the cycle time of the injection molding machine from four minutes to three minutes per part. Thirty percent of the direct labor and 45% of the factory overhead are related to running injection molding machines.

The direct labor rate is $21 per hour.

a. Determine the target cost for Model J20 assuming that the historical markup on product cost and selling price are maintained. Round your final answer to two decimal places.
$

b. Determine the required cost reduction. Enter as a positive number. Round your final answer to two decimal places.
$

c. Evaluate the three engineering improvements together to determine if the required cost reduction (drift) can be achieved. Enter all amounts as positive numbers. Do not round interim calculations but round your final answers to two decimal places.

1. Direct labor reduction $
2. Additional inspection $
3. Injection molding productivity improvement $
Total savings $
Answer #1
Markup % 25.00%
New Selling Price 330
Target Cost 247.50
Required cost reduction 32.50
1. Direct labor reduction 3.15
2. Additional inspection 4.9
3. Injection molding productivity improvement 21.13875
Total savings 29.18875
No, will not be achieved

Α. 1 Markup % =(350-280)/280 2 3 New Selling Price 4 Target Cost 330 =B3*(1-B1) 6 Required cost reduction =280-B4 7 8 1. Dire